OpenStudy (anonymous):

3. Michelle invests $8000 in a money market account that pays 3.25% interest compounded monthly. How much will be in the account after 12 years? Example:

OpenStudy (anonymous):

Okay. This question is almost exactly like the last one you posted. There's one VERY important difference now: This question tells us that the interest increases 3.25% every MONTH. NOT every year. So, we have to be careful. ~~~~~~~~~~ Here, again, is the equation: \[y=a (1+b)^{x}\] Y is the answer to your question. A is the starting amount of money. (In this case, it would be 8000.) B is the percentage of increase, in decimal form. (In this case, 3.25% would be 0.0325) X is the number of MONTHS passed. We have to use MONTHS, because the interest increases MONTHLY (as the problem says. 12 years = 24 months. ~~~~~~~~~~~ So, once again, why don't you try plugging in the numbers. When you have an answer, let me know. If you have trouble, let me know.

OpenStudy (anonymous):

y=8000(1+0.0335)^24

OpenStudy (anonymous):

12 yrs is 144 months everyone... 12 months a year times 12 years is 144 months... so the equation should read y=8000(1+0.0335)^144=920024.18

OpenStudy (anonymous):

@Mr.Turner @kingdaddy100 That's right, 12 years is 144 months. I'm sorry, for some reason, I read 2 years, instead of 12.

OpenStudy (anonymous):

i thought it was 3.25% not 3.35% so the equation is y=8000(1+0.0325)^144=920024.18 so y=800(1.0325)^144 y=80 (100.0388) y= 8003.104 y=800310.7938

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

The increase is 3.25%. That's what the problem tells us. So, the decimal would be .0325 Your equation would be... y = 8000(1.0325)^144 The parentheses is already taken care of. We need to take care of the exponent first, and then multiply. In the end, we would have $800,310.79

OpenStudy (anonymous):

thanks